# Incentivizing Exploration with Linear Contexts and Combinatorial Actions

### v2

@@ Abstract
We advance the study of incentivized bandit exploration, in which arm choices are viewed as recommendations and are required to be Bayesian incentive compatible. Recent work has shown under certain independence assumptions that after collecting enough initial samples, the popular Thompson sampling algorithm becomes incentive compatible. We give an analog of this result for linear bandits, where the independence of the prior is replaced by a natural convexity condition. This opens up the possibility of efficient and regret-optimal incentivized exploration in high-dimensional action spaces. In the semibandit model, we also improve the sample complexity for the pre-Thompson sampling phase of initial data collection.
@@ 1.1 Problem Formulation and Results
Let ${\mathcal{A}}$ be an action set and $\mu$ a prior distribution over the mean reward function *$\ell^{*}:{\mathcal{A}}\to{\mathbb{R}}$. A bandit algorithm which recommends a sequence $(A^{(1)},A^{(2)},\dots,A^{(T)})$ of actions in ${\mathcal{A}}$ is said to be Bayesian incentive compatible (BIC) if for each $t\in[T]$:
@@
As mentioned, fundamental results on BIC bandit algorithms were obtained in [MSS20, MSSW22]. For instance, the former work gave “hidden exploration” algorithms which exploit with such high frequency that performing any bandit algorithm on the remaining time-steps is BIC (given suitable assumptions on the prior). Unfortunately, all algorithms in these works had to pay exponentially large multiplicative factors in their regret compared to non-BIC bandit algorithms. This “price of incentives” was studied quantitatively in our recent work [SS23] for the case of independent arms. They proved the classical Thompson sampling algorithm [Tho33] is BIC without modification once a mild number of initial samples per arm have been collected, leading to a natural two-phase algorithm. The first phase focused on collecting these initial data in a BIC way, while the second is simply Thompson sampling. Note that Thompson sampling obeys many state-of-the-art regret bounds, both Bayesian [RV14, BL13] and frequentist [AG12, KKM12, AG17]. Thus the result of [SS23] implies that the additional regret forced by the BIC requirement is essentially additive, and at most the number of rounds needed to collect the needed initial samples in a BIC way (a quantity studied separately therein).
@@
The results of [SS23] are limited to the case of independent arms, and it is therefore of interest to expand the range of models under which Thompson sampling leads to provable incentive compatibility. This is the goal of the present paper. Our main focus is on the linear bandit in dimension $d$. Here the action set ${\mathcal{A}}\subseteq{\mathbb{R}}^{d}$ may be infinite, and the reward function *$\ell^{*}:{\mathcal{A}}\to{\mathbb{R}}$ is always linear. This is a natural next step and allows for richer correlations between actions. From the practical viewpoint, BIC guarantees are relevant for recommending restaurants, movies, hotels, and doctors (see e.g. [Sli19], Chapter 11); in all of these settings it is desirable to leverage contextual information to obtain sample complexity scaling with the ambient data dimension rather than the potentially huge number of total actions. We make no assumptions of independence but instead require natural geometric conditions, e.g. that the prior $\mu$ for $\ell^{*}$ is uniformly random on a convex body ${\mathcal{K}}$ with bounded aspect ratio. This covers a fairly broad range of scenarios, including centered Gaussian $\mu$ by homogeneity. Indeed we consider the connection with such conditions as a significant contribution of this work; in [SS23, HNSW22] the FKG inequality is used crucially throughout but requires independence properties that are unavaiable with linear contexts. As a regret benchmark, recall that for the Bayesian regret of Thompson sampling is known to be $O(d\sqrt{T\log T})$ by [DVR18, Theorem 2] (see also [RV14]).
@@
Relative to this near-optimal guarantee, our first main result Theorem 3.5 shows that the price of incentives is again additive rather than multiplicative. Namely, Thompson sampling is again BIC after obtaining ${\rm poly}(d)$ initial samples which are well-spread in a natural spectral sense. Technically, we require action sets be to $\varepsilon$-separated so that the recommendation of a given $A\in{\mathcal{A}}$ has non-negligible probability; this is a mild condition since any action set can be discretized before running Thompson sampling. Further as shown in Theorem 3.7, the result extends to the generalized linear bandit when the link function’s derivative is bounded above and below. We finally give two counterexamples. The former shows that Thompson sampling may be BIC at time $1$ but not time $2$ – this may be surprising as it is impossible for the multi-armed bandit with independent arms (see [SS23, Lemma 4.9]). The latter gives a natural example in which initial data collection provably needs $e^{\Omega(d)}$ samples. In particular, the latter counterexample illustrates that further geometric conditions on the prior and/or action set are necessary for ${\rm poly}(d)$ sample complexity of BIC initial exploration. We leave a further study of this interesting problem for future work, but remark that exogenous payments could be used in practice to obtain the required ${\rm poly}(d)$ initial samples (see some of the references in Subsection 1.2).
@@ 1.2 Other Relevant Work
Here we mention a few other related works in the broad area of incentivized exploration. [FKKK14] studies a similar problem, again in a Bayesian setting. However in their work the full history is made public so there is no information asymmetry. Instead, incentivization is achieved via exogenous payments. [WH18, AT20, ${}^{+}$23] study incentivized exploration in a non-Bayesian setting, where empirical averages are used instead of posterior means, and again incentives are realized through payments. The latter two works also focus on linear contexts and show $\widetilde{O}(\sqrt{T})$ total payment suffices for incentivization. By contrast Theorem 3.5 implies that Thompson sampling is BIC after a constant (i.e. $T$-independent) amount of initial exploration. If payments can be used for incentivization, this implies in particular that a constant amount of total payment suffices for incentivized exploration. However as just mentioned, our setting and assumptions differ from the aforementioned works in multiple ways.
[${}^{+}$17] studied the power of exogenous payments to incentivize related notions of fairness in which better actions ought to be played with higher probability. Their model includes information asymmetry in the opposite direction; agents observe the full history while the planner might not. [IMSW20] proposed incentivization via selective data disclosure where agents observe a carefully chosen subset of the history. However their setting is not precisely Bayesian as they assume agents perform naive empirical averaging over the chosen subset rather than taking the planner’s algorithm into account. In fact it can be shown by a form of the revelation princple (see [Sli19], Chapter 11) that in the perfectly Bayesian setting, general planner-to-agent signals have no more power than simple action recommendations. Finally our work and the ones mentioned above assume agents are identical; a model with heterogenous agents was studied in [IMSW19].
@@ 2 Preliminaries
where $z_{t}$ is conditionally mean zero and $O(1)$-subgaussian conditioned on (*$(\ell^{*},A^{(t)})$. We assume $\ell^{*}$ is such that
@@ 3 Linear Bandit
If ${\mathcal{A}}=(e_{1},\dots,e_{d})$ forms an orthonormal basis (as in the multi-armed bandit setting), this simply means that each action was sampled at least $\gamma$ times. We say that $\gamma$-spectral exploration has occurred at time $t$ if (3.1) holds. Note that if one is willing to simply purchase initial samples, then $\gamma d$ samples are typically needed to achieve $\gamma$-spectral exploration. The next lemma follows from Lemma 2.2 and is proved in the appendix. We note that the factor of $\sqrt{d\log(t)}$ comes from the adaptivity of the exploration, as explained in [LS20, Exercise 20.2]. We thank Anand Kalvit for bringing this to our attention, and thus correcting an error in an earlier version of this paper.
@@
Suppose $\gamma$-spectral exploration has occurred almost surely at some (deterministic) time $t$. Then the random vector *$\ell^{*}-{\mathbb{E}}^{t}[\ell^{*}]$ has zero mean and is $O(\sqrt{d\log(t)/\gamma})$-subgaussian.
@@ 3.1 Convex Geometry
Suppose that ${\mathcal{A}}\subseteq{\mathbb{S}}^{d-1}$ is $\varepsilon$-separated, the convex set ${\mathcal{K}}\subseteq{\mathbb{R}}^{d}$ is $r$-regular, and *$\ell^{*}\sim\mu={\cal U}({\mathcal{K}})$. Then for each $A_{i},A_{j}\in{\mathcal{A}}$, we have
@@ 3.2 Main Result for Linear Bandit
There exists an absolute constant $C$ such that the following holds. Suppose that ${\mathcal{A}}\subseteq{\mathbb{S}}^{d-1}$ is $\varepsilon$-separated, the convex set ${\mathcal{K}}\subseteq{\mathbb{R}}^{d}$ is $r$-regular, and *$\ell^{*}\sim\mu={\cal U}({\mathcal{K}})$. If (3.1) holds almost surely at time $t$ for
@@ 3.3 Extension to Generalized Linear Bandit
for a known strictly increasing link function $\chi:{\mathbb{R}}\to{\mathbb{R}}$ and unknown vector $\ell^{*}$. A notable example is the logistic bandit where $\chi(x)=\frac{e^{x}}{1+e^{x}}$. Recalling that we assumed *$\langle\ell^{*},x\rangle\in[-1,1]$ for all *$\ell^{*}\in{\rm supp}(\mu)$ and $x\in A$, we follow much of the theoretical literature on this model by fixing the non-negative constants
@@
Fix a link function $\mu:{\mathbb{R}}\to{\mathbb{R}}$. Suppose that ${\mathcal{A}}\subseteq{\mathbb{S}}^{d-1}$ is $\varepsilon$-separated, the convex set ${\mathcal{K}}\subseteq{\mathbb{R}}^{d}$ is $r$-regular, and *$\ell^{*}\sim\mu={\cal U}({\mathcal{K}})$, where $r,\varepsilon,m_{\chi}\geq e^{-d}$. Given deterministic initial exploration until time $t$ under which (3.1) holds for
@@ 3.4 Two Counterexamples
or its set of extreme points. We show that for a suitable prior $\mu$, any initial exploration algorithm which is BIC or even $\varepsilon$-BIC can be made to require $e^{\Omega(d)}$ timesteps to explore in the $d$-th coordinate.
Note that for ${\mathcal{P}}$ as above, all extreme points satisfy $x_{d}=0$ except for ±$\pm\hat{A}=\left(0,0,\dots,0,\pm\frac{1}{2\sqrt{d}}\right)$, and so $\gamma$-spectral exploration for any $\gamma>0$ requires exploring these actions. Our prior distribution *$\ell^{*}\sim\mu$ will be uniform over the “biased” convex body ${\mathcal{K}}/\sqrt{d}$ for
@@
With ${\mathcal{A}},\mu$ as above, suppose that an $\varepsilon$-BIC algorithm ${\mathbb{A}}$ has explored at least one of $\hat{A}$ or $-\hat{A}$ almost surely by time $T$ for $\varepsilon\leq 1/20$. Then $T\geq\exp(\Omega(d))$.
@@ A.1 Proofs for Generalized Linear Bandit
[|^*${\mathbb{P}}\left[|\langle\hat{\ell}-\ell^{*},v\rangle|\leq\frac{\|v\|}{m_{\chi}\sqrt{\gamma}}\right]\geq 1-\delta.$
@@
Fix a link function $\chi$. Suppose that ${\mathcal{A}}\subseteq{\mathbb{S}}^{d-1}$ is $\varepsilon$-separated, the convex set ${\mathcal{K}}\subseteq{\mathbb{R}}^{d}$ is $r$-regular, and *$\ell^{*}\sim\mu={\cal U}({\mathcal{K}})$. Then for each $A_{i},A_{j}\in{\mathcal{A}}$, we have
@@ B.1 Exponential Lower Bound for Initial Exploration
Suppose there exists a BIC algorithm ${\mathcal{A}}$ which explores in ${\mathcal{P}}$ and achieves $\gamma$-spectral exploration almost surely within $T$ timesteps. Then there exists a BIC algorithm ${\mathcal{A}}^{\prime}$ which explores in ${\mathcal{A}}={\mathcal{E}}({\mathcal{P}})$ and achieves $\gamma$-spectral exploration almost surely within $Td$ timesteps.
@@
We show how to simulate each step of ${\mathcal{A}}$ using exactly $d$ steps of ${\mathcal{A}}^{\prime}$, in a BIC way. For each $t\in\{0,1,\dots,T-1\}$, at the start of time-step $td+1$ we will have a simulated version of ${\mathcal{A}}$ which has received some artifical feedback and recommends an action $A^{(t)}\in{\mathcal{P}}$. To construct ${\mathcal{A}}^{\prime}$ we do the following:
@@
It is easy to see that $r_{t}$ has the correct expected value by linearity of rewards for the linear bandit problem. It then follows that after receiving $(r_{1},\dots,r_{t-1})$, the recommendation $A^{(t)}\in{\mathcal{P}}$ is BIC within ${\mathcal{P}}$ by the BIC property of ${\mathcal{A}}$. By linearity it follows that conditioned on observing $A^{(t)}$, each extreme point $\widetilde{A}_{t,i}$ is also a BIC recommendation, since they all must have the same conditional mean reward.
It remains to show that ${\mathcal{A}}^{\prime}$ also achieves $\gamma$-spectral exploration. Indeed
@@
Based on Proposition B.1, we state Proposition 3.9 below in the discrete action set setting ${\mathcal{A}}={\mathcal{E}}({\mathcal{P}})$, with the understanding that it extends also to the convex action set ${\mathcal{A}}={\mathcal{P}}$. Note that for ${\mathcal{P}}$ as above, all extreme points satisfy $x_{d}=0$ except for ±$\pm\hat{A}=\left(0,0,\dots,0,\pm\frac{1}{2\sqrt{d}}\right)$, and so $\gamma$-spectral exploration for any $\gamma>0$ requires exploring these actions. Our prior distribution *$\ell^{*}\sim\mu$ will be uniform over the “biased” convex body ${\mathcal{K}}/\sqrt{d}$ for
@@
Consider a modification ${\mathbb{A}}^{\prime}$ of ${\mathbb{A}}$ which changes all plays of ±$\pm\hat{A}$ to the prior-optimal action $A_{1}=\frac{(1,1,\dots,1,0)}{2\sqrt{d}}$. If ${\mathbb{A}}^{\prime}$ and ${\mathbb{A}}$ play different actions at $m$ times in expectation, then by definition the expected total rewards satisfy
$R_{T}({\mathbb{A}}^{\prime})+\varepsilon m\geq R_{T}({\mathbb{A}}).$ (B.2)
@@
${\mathbb{E}}[r_{\tau}({\mathbb{A}}^{\prime})+\varepsilon-r_{\tau}({\mathbb{A}})]$ $\leq\varepsilon-{\mathbb{E}}[\langle\ell^{*},A_{1}\rangle]$
$\leq\varepsilon-1/3\leq-1/6.$
@@
${\mathbb{E}}^{s}[r_{s}({\mathbb{A}}^{\prime})+\varepsilon-r_{s}({\mathbb{A}})]\geq\Big{(}|{\mathbb{E}}^{s}[\langle\ell^{*},\hat{A}\rangle]|+\varepsilon-{\mathbb{E}}^{s}[\langle\ell^{*},A_{1}\rangle]\Big{)}_{+}.$
@@
$R_{T}({\mathbb{A}}^{\prime})+\varepsilon m-R_{T}({\mathbb{A}})$ $\leq-1/6+\sum_{s=1}^{T}{\mathbb{E}}\left[\Big{(}|{\mathbb{E}}^{s}[\langle\ell^{*},\hat{A}\rangle]|+\varepsilon-{\mathbb{E}}^{s}[\langle\ell^{*},A_{1}\rangle]\Big{)}_{+}\leavevmode\nobreak\ \right]$
$\leq-1/6+T\cdot{\mathbb{E}}\left[\Big{(}|\langle\ell^{*},\hat{A}\rangle|+\varepsilon-\langle\ell^{*},A_{1}\rangle\Big{)}_{+}\leavevmode\nobreak\ \right].$

### v3

@@ Abstract
We advance the study of incentivized bandit exploration, in which arm choices are viewed as recommendations and are required to be Bayesian incentive compatible. Recently [SS23] showed under certain independence assumptions that after collecting enough initial samples, the popular Thompson sampling algorithm becomes incentive compatible. We give an analog of this result for linear bandits, where the independence of the prior is replaced by a natural convexity condition. This opens up the possibility of efficient and regret-optimal incentivized exploration in high-dimensional action spaces. In the semibandit model, we also improve the sample complexity for the pre-Thompson sampling phase of initial data collection.
@@ 1.1 Problem Formulation and Results
Let ${\mathcal{A}}$ be an action set and $\mu$ a prior distribution over the mean reward function $\ell^{*}:{\mathcal{A}}\to{\mathbb{R}}$. A bandit algorithm which recommends a sequence $(A^{(1)},A^{(2)},\dots,A^{(T)})$ of actions in ${\mathcal{A}}$ is said to be Bayesian incentive compatible (BIC) if for each $t\in[T]$:
@@
As mentioned, fundamental results on BIC bandit algorithms were obtained in [MSS20, MSSW22]. For instance, the former work gave “hidden exploration” algorithms which exploit with such high frequency that performing any bandit algorithm on the remaining time-steps is BIC (given suitable assumptions on the prior). Unfortunately, all algorithms in these works had to pay exponentially large multiplicative factors in their regret compared to non-BIC bandit algorithms. This “price of incentives” was studied quantitatively in our recent work [SS23] with Slivkins for the case of independent arms. We proved the classical Thompson sampling algorithm [Tho33] is BIC without modification once a mild number of initial samples per arm have been collected, leading to a natural two-phase algorithm. The first phase collects these initial samples in a BIC way, while the second simply performs Thompson sampling. Note that Thompson sampling obeys many state-of-the-art regret bounds, both Bayesian [RV14, BL13, BS22] and frequentist [AG12, KKM12, AG17, ZL19, LG21]. Thus the result of [SS23] implies that the additional regret forced by the BIC requirement is essentially additive, and at most the number of rounds needed to collect the needed initial samples in a BIC way (a quantity studied separately therein).
@@
The results of [SS23] are limited to the case of independent arms, and it is therefore of interest to expand the range of models under which Thompson sampling leads to provable incentive compatibility. This is the goal of the present paper. Our main focus is on the linear bandit in dimension $d$. Here the action set ${\mathcal{A}}\subseteq{\mathbb{R}}^{d}$ may be infinite, and the reward function $\ell^{*}:{\mathcal{A}}\to{\mathbb{R}}$ is always linear. This is a natural next step and allows for richer correlations between actions. From the practical viewpoint, BIC guarantees are relevant for recommending restaurants, movies, hotels, and doctors (see e.g. [Sli19], Chapter 11); in all of these settings it is desirable to leverage contextual information to obtain sample complexity scaling with the ambient data dimension rather than the potentially huge number of total actions. We make no assumptions of independence but instead require natural geometric conditions, e.g. that the prior $\mu$ for $\ell^{*}$ is uniformly random on a convex body ${\mathcal{K}}$ with bounded aspect ratio. This covers a fairly broad range of scenarios, including centered Gaussian $\mu$ by homogeneity. Indeed we consider the connection with such conditions as a significant contribution of this work; in [SS23, HNSW22] the FKG inequality is used crucially throughout but requires independence properties that are unavailable with linear contexts. As a regret benchmark, recall that for the Bayesian regret of Thompson sampling is known to be $O(d\sqrt{T\log T})$ by [DV18, Theorem 2] (see also [RV14]). Interestingly, the frequentist regret of Thompson sampling for linear bandits is known to be larger by a factor of $\sqrt{d}$ [AG13, HB20].
@@
Relative to this near-optimal guarantee, our first main result Theorem 3.5 shows that the price of incentives is again additive rather than multiplicative. Namely, Thompson sampling is again BIC after obtaining ${\rm poly}(d)$ initial samples which are well-spread in a natural spectral sense. Technically, we require action sets be to $\varepsilon$-separated so that the recommendation of a given $A\in{\mathcal{A}}$ has non-negligible probability; this is a mild condition since any action set can be discretized before running Thompson sampling. Further as shown in Theorem 3.7, the result extends to the generalized linear bandit when the link function’s derivative is bounded above and below.
Next, we provide two counterexamples. The former shows that Thompson sampling may be BIC at time $1$ but not time $2$ – this may be surprising as it is impossible for the multi-armed bandit with independent arms (see [SS23, Lemma 4.9]). The latter gives a natural example in which initial data collection provably needs $e^{\Omega(d)}$ samples. In particular, the latter counterexample illustrates that further geometric conditions on the prior and/or action set are necessary for ${\rm poly}(d)$ sample complexity of BIC initial exploration. We leave a further study of this interesting problem for future work, but remark that exogenous payments could be used in practice to obtain the required ${\rm poly}(d)$ initial samples (see some of the references in Subsection 1.2).
@@ 1.2 Other Relevant Work
Here we mention a few other related works in the broad area of incentivized exploration. [FKKK14] studies a similar problem, again in a Bayesian setting. However in their work the full history is made public so there is no information asymmetry. Instead, incentivization is achieved via exogenous payments. [WH18, AT20, WXL+23] study incentivized exploration in a non-Bayesian setting, where empirical averages are used instead of posterior means, and again incentives are realized through payments. The latter two works also focus on linear contexts and show $\widetilde{O}(\sqrt{T})$ total payment suffices for incentivization. By contrast Theorem 3.5 implies that Thompson sampling is BIC after a constant (i.e. $T$-independent) amount of initial exploration. If payments can be used for incentivization, this implies in particular that a constant amount of total payment suffices for incentivized exploration. However as just mentioned, our setting and assumptions differ from the aforementioned works in multiple ways.
[KKM+17] studied the power of exogenous payments to incentivize related notions of fairness in which better actions ought to be played with higher probability. Their model includes information asymmetry in the opposite direction; agents observe the full history while the planner might not. [IMSW20] proposed incentivization via selective data disclosure where agents observe a carefully chosen subset of the history. However their setting is not precisely Bayesian as they assume agents perform naive empirical averaging over the chosen subset rather than taking the planner’s algorithm into account. In fact it can be shown by a form of the revelation princple (see [Sli19], Chapter 11) that in the perfectly Bayesian setting, general planner-to-agent signals have no more power than simple action recommendations. Finally our work and the ones mentioned above assume agents are identical; a model with heterogenous agents was studied in [IMSW19].
@@ 2 Preliminaries
where $z_{t}$ is conditionally mean zero and $O(1)$-subgaussian conditioned on $(\ell^{*},A^{(t)})$. We assume $\ell^{*}$ is such that
@@ 3 Linear Bandit
If ${\mathcal{A}}=(e_{1},\dots,e_{d})$ forms an orthonormal basis (as in the multi-armed bandit setting), this simply means that each action was sampled at least $\gamma$ times. We say that $\gamma$-spectral exploration has occurred at time $t$ if (3.1) holds. Note that if one is willing to simply purchase initial samples, then $\gamma d$ samples are typically needed to achieve $\gamma$-spectral exploration. The next lemma follows from Lemma 2.2 and is proved in the appendix. We note that the factor of $\sqrt{d\log(t)}$ comes from the adaptivity of the exploration, as explained in [LS20, Exercise 20.2].
@@
Suppose $\gamma$-spectral exploration has occurred almost surely at some (deterministic) time $t$. Then the random vector $\ell^{*}-{\mathbb{E}}^{t}[\ell^{*}]$ has zero mean and is $O(\sqrt{d\log(t)/\gamma})$-subgaussian.
@@ 3.1 Convex Geometry
Suppose that ${\mathcal{A}}\subseteq{\mathbb{S}}^{d-1}$ is $\varepsilon$-separated, the convex set ${\mathcal{K}}\subseteq{\mathbb{R}}^{d}$ is $r$-regular, and $\ell^{*}\sim\mu={\cal U}({\mathcal{K}})$. Then for each $A_{i},A_{j}\in{\mathcal{A}}$, we have
@@ 3.2 Main Result for Linear Bandit
There exists an absolute constant $C$ such that the following holds. Suppose that ${\mathcal{A}}\subseteq{\mathbb{S}}^{d-1}$ is $\varepsilon$-separated, the convex set ${\mathcal{K}}\subseteq{\mathbb{R}}^{d}$ is $r$-regular, and $\ell^{*}\sim\mu={\cal U}({\mathcal{K}})$. If (3.1) holds almost surely at time $t$ for
@@ 3.3 Extension to Generalized Linear Bandit
for a known strictly increasing link function $\chi:{\mathbb{R}}\to{\mathbb{R}}$ and unknown vector $\ell^{*}$. A notable example is the logistic bandit where $\chi(x)=\frac{e^{x}}{1+e^{x}}$. Recalling that we assumed $\langle\ell^{*},x\rangle\in[-1,1]$ for all $\ell^{*}\in{\rm supp}(\mu)$ and $x\in A$, we follow much of the theoretical literature on this model by fixing the non-negative constants
@@
Fix a link function $\mu:{\mathbb{R}}\to{\mathbb{R}}$. Suppose that ${\mathcal{A}}\subseteq{\mathbb{S}}^{d-1}$ is $\varepsilon$-separated, the convex set ${\mathcal{K}}\subseteq{\mathbb{R}}^{d}$ is $r$-regular, and $\ell^{*}\sim\mu={\cal U}({\mathcal{K}})$, where $r,\varepsilon,m_{\chi}\geq e^{-d}$. Given deterministic initial exploration until time $t$ under which (3.1) holds for
@@ 3.4 Two Counterexamples
or its set of extreme points ${\mathcal{E}}({\mathcal{P}})$. We show that for a suitable prior $\mu$, any initial exploration algorithm which is BIC or even $\varepsilon$-BIC can be made to require $e^{\Omega(d)}$ timesteps to explore in the $d$-th coordinate.
Note that for ${\mathcal{P}}$ as above, all extreme points satisfy $x_{d}=0$ except for ±$\pm\hat{A}=\left(0,0,\dots,0,\pm\frac{1}{10}\right)$, and so $\gamma$-spectral exploration for any $\gamma>0$ requires exploring these actions. Our prior distribution $\ell^{*}\sim\mu$ will be uniform over the “biased” convex body ${\mathcal{K}}/\sqrt{d}$ for
@@
With ${\mathcal{A}},\mu$ as above for $d\geq 100$, suppose that ${\mathbb{A}}$ is either an $0.01$-BIC algorithm on action set ${\mathcal{E}}({\mathcal{P}})$, or a BIC algorithm on ${\mathcal{P}})$. If at least one of $\hat{A}$ or $-\hat{A}$ is almost surely explored by ${\mathbb{A}}$ before time $T$, then $T\geq\exp(\Omega(d))$.
@@ 4.3 Lower Bound
Acknowledgement
Thanks to Anand Kalvit for bringing [LS20, Exercise 20.2] to our attention, and thus correcting an error in the regret bounds for linear bandits. Thanks also to Ben Schiffer for pointing out a number of typos in the second counterexample.
@@ A.1 Proofs for Generalized Linear Bandit
[|^${\mathbb{P}}\left[|\langle\hat{\ell}-\ell^{*},v\rangle|\leq\frac{\|v\|}{m_{\chi}\sqrt{\gamma}}\right]\geq 1-\delta.$
@@
Fix a link function $\chi$. Suppose that ${\mathcal{A}}\subseteq{\mathbb{S}}^{d-1}$ is $\varepsilon$-separated, the convex set ${\mathcal{K}}\subseteq{\mathbb{R}}^{d}$ is $r$-regular, and $\ell^{*}\sim\mu={\cal U}({\mathcal{K}})$. Then for each $A_{i},A_{j}\in{\mathcal{A}}$, we have
@@ B.1 Exponential Lower Bound for Initial Exploration
Suppose there exists a BIC algorithm ${\mathbb{A}}$ which explores in ${\mathcal{P}}$ and achieves $\gamma$-spectral exploration almost surely within $T$ timesteps. Then there exists a BIC algorithm ${\mathbb{A}}^{\prime}$ which explores in ${\mathcal{A}}={\mathcal{E}}({\mathcal{P}})$ and achieves $\gamma$-spectral exploration almost surely within $Td$ timesteps.
@@
We show how to simulate each step of ${\mathbb{A}}$ using exactly $d$ steps of ${\mathbb{A}}^{\prime}$, in a BIC way. For each $t\in\{0,1,\dots,T-1\}$, at the start of time-step $td+1$ we will have a simulated version of ${\mathbb{A}}$ which has received some artifical feedback and recommends an action $A^{(t)}\in{\mathcal{P}}$. To construct ${\mathbb{A}}^{\prime}$ we do the following:
@@
It is easy to see that $r_{t}$ has the correct expected value by linearity of rewards for the linear bandit problem. It then follows that after receiving $(r_{1},\dots,r_{t-1})$, the recommendation $A^{(t)}\in{\mathcal{P}}$ is BIC within ${\mathcal{P}}$ by the BIC property of ${\mathbb{A}}$. By linearity it follows that conditioned on observing $A^{(t)}$, each extreme point $\widetilde{A}_{t,i}$ is also a BIC recommendation, since they all must have the same conditional mean reward.
It remains to show that ${\mathbb{A}}^{\prime}$ also achieves $\gamma$-spectral exploration. Indeed
@@
Based on Proposition B.1, we state Proposition 3.9 below in the discrete action set setting ${\mathcal{A}}={\mathcal{E}}({\mathcal{P}})$ and an assumption of $0.01$-BIC, with the understanding that it extends also to the convex action set ${\mathcal{A}}={\mathcal{P}}$ for BIC algorithms. Note that for ${\mathcal{P}}$ as above, all extreme points satisfy $x_{d}=0$ except for ±$\pm\hat{A}=\left(0,0,\dots,0,\pm\frac{1}{10}\right)$, and so $\gamma$-spectral exploration for any $\gamma>0$ requires exploring these actions. Our prior distribution $\ell^{*}\sim\mu$ will be uniform over the “biased” convex body ${\mathcal{K}}/\sqrt{d}$ for
@@
We set $\varepsilon=0.01$. Consider a modification ${\mathbb{A}}^{\prime}$ of ${\mathbb{A}}$ which changes all plays of ±$\pm\hat{A}$ to the prior-optimal action $A_{1}=\frac{(1,1,\dots,1,0)}{2\sqrt{d}}$. If ${\mathbb{A}}^{\prime}$ and ${\mathbb{A}}$ play different actions at $m$ times in expectation, then by definition the expected total rewards satisfy
$R_{T}({\mathbb{A}})+\varepsilon m\geq R_{T}({\mathbb{A}}^{\prime}).$ (B.2)
@@
${\mathbb{E}}[r_{\tau}({\mathbb{A}})+\varepsilon-r_{\tau}({\mathbb{A}}^{\prime})]\leq\varepsilon-{\mathbb{E}}[\langle\ell^{*},A_{1}\rangle]\leq\varepsilon-\frac{d-1}{8d}\leq-1/9.$
@@
${\mathbb{E}}^{s}[r_{s}({\mathbb{A}})+\varepsilon-r_{s}({\mathbb{A}}^{\prime})]\leq\Big{(}|{\mathbb{E}}^{s}[\langle\ell^{*},\hat{A}\rangle]|+\varepsilon-{\mathbb{E}}^{s}[\langle\ell^{*},A_{1}\rangle]\Big{)}_{+}.$
@@
$R_{T}({\mathbb{A}})+\varepsilon m-R_{T}({\mathbb{A}}^{\prime})$ $\leq-1/9+\sum_{s=1}^{T}{\mathbb{E}}\left[\Big{(}|{\mathbb{E}}^{s}[\langle\ell^{*},\hat{A}\rangle]|+\varepsilon-{\mathbb{E}}^{s}[\langle\ell^{*},A_{1}\rangle]\Big{)}_{+}\leavevmode\nobreak\ \right]$
$\leq-1/9+T\cdot{\mathbb{E}}\left[\Big{(}|\langle\ell^{*},\hat{A}\rangle|+\varepsilon-\langle\ell^{*},A_{1}\rangle\Big{)}_{+}\leavevmode\nobreak\ \right].$